Morganton Company makes one product and it provided the following information to help prepare the master budget:

  1. The budgeted selling price per unit is $70. Budgeted unit sales for June, July, August, and September are 8,400, 10,000, 12,000, and 13,000 units, respectively. All sales are on credit.
  2. Forty percent of credit sales are collected in the month of the sale and 60% in the following month.
  3. The ending finished goods inventory equals 20% of the following month’s unit sales.
  4. The ending raw materials inventory equals 10% of the following month’s raw materials production needs. Each unit of finished goods requires 5 pounds of raw materials. The raw materials cost $2.00 per pound.
  5. Thirty percent of raw materials purchases are paid for in the month of purchase and 70% in the following month.
  6. The direct labor wage rate is $15 per hour. Each unit of finished goods requires two direct labor-hours.
  7. The variable selling and administrative expense per unit sold is $1.80. The fixed selling and administrative expense per month is $60,000.

15. If we assume that there is no fixed manufacturing overhead and the variable manufacturing overhead is $10 per direct labor-hour, what is the estimated net operating income for July?

Budgeted Sales : $700,000

Total Cash Collections : $632,800

Accounts Receivable : $420,000

Required Production : 10,400 units

Raw Materials To Be Purchased : 52,900 pounds

Cost Of Raw Materials To Be Purchased : $105,800

Total Cash Disbursements : $93,956

Accounts Payable : $74,060

Raw Materials Inventory Balance : $12,200
Answer #1
Direct materials 10 =5*2
Direct labor 30 =15*2
Variable manufacturing overhead 20 =10*2
Unit product cost 60
15
Sales 700000 =10000*70
Less: Cost of goods sold 600000 =10000*60
Gross margin 100000
Less: Selling and administrative expense 78000 =60000+(10000*1.80)
Estimated net operating income for July 22000